A particle leaves the 'origin with velocity u. If
it moves with an acceleration =m/v^2,where v be
the velocity at any instant, show that the
distance x travelled in time t is given by
4 cx= (3 ct+ c^2)4/3 - u^4.
1
Expert's answer
2020-10-09T07:13:43-0400

As per the question,

Initial velocity of the particle is u,

Acceleration of the particle "(a)=\frac{m}{v^2}"

Let the distance traveled by the particle in the time t is x

"a=\frac{dv}{dt}"

"\Rightarrow dv= a dt"

Now, substituting the values of a in the above,

"dv=\frac{m}{v^2}dt"

Now, taking the integration,

"\int_u^v dv=\int_0^t\frac{mdt}{v^2}"


"\Rightarrow \int_u^v v^2dv=\int_0^t mdt"


"\Rightarrow [\frac{v^3}{3}]_u^v=mt"


"\Rightarrow \frac{v^3-u^3}{3}=mt"

"v=(3mt+u^3)^{1\/3}"

Now, we know that "v= \frac{dx}{dt}"

"dx=vdt"

Now, taking the integration of both side,

"\int_0^xdx=\int_0^tvdt"

Now , substituting the values,

"x=\int_0^t(3mt+u^3)^{1/3})dt"

"\Rightarrow x=[\frac{(3mt+u^3)^{4/3}}{4m }]_0^t"

"\Rightarrow x=[\frac{(3mt+u^3)^{4/3}-u^4}{4m }]"

"\Rightarrow 4mx=(3mt+u^3)^{4/3}-u^4"